OpenStudy (anonymous):

find dy/dx for 2y^3+ xy+1=0

OpenStudy (anonymous):

\[6y^2y \prime+y+xy \prime=0\]

OpenStudy (anonymous):

\[xy=(x)^{'}y+(y )^{'}x\] product rule

OpenStudy (anonymous):

Simplest way first make x a Dependant variable, and then find dx/dy. The resprocal of it will be the answer

OpenStudy (anonymous):

is dy/dx of 1 = 0?

OpenStudy (anonymous):

yes dy/dx of 1 does =0

OpenStudy (anonymous):

\[6y^2y \prime + xy \prime + y\]

OpenStudy (anonymous):

\[y^{'}=\frac{ -y }{ 6y^2+x }\] implicit differentiation

OpenStudy (anonymous):

\[y ^{'}=dy/dx\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

.